OpenStudy (anonymous):

Select Yes or No for each question. Do the following sets of numbers create a triangle? yes or no: 900 cm, 60 cm, 0.5 m yes or no: 12 m, 34 m, 40 m yes or no: 145 ft, 10 yd, 105 ft

OpenStudy (anonymous):

no yes no

OpenStudy (anonymous):

thanks(:

OpenStudy (anonymous):

wels

OpenStudy (anonymous):

Basic rule for a triangle is that the sum of any two sides must be greater than the other side. 60+0.5 < 900 = not a triangle. 12+34 > 40, 12+40 > 34, 34+40 > 12 = triangle 105+30 < 145 = not triangle